# This is a preprint submitted to: Journal of Hydrodynamics

### v1

@@ 1 Test Case
A fiberglass scale model, designated as CNR-INM 2554, was constructed to replicate the Delft design at a scale of 1:33.33, corresponding to a model length (Lpp) of 3 meters, considering the full-scale prototype to be a notional 100-meter catamaran. An outline of the model’s geometry and body plan is illustrated in Fig.˜1, with a visual of the manufactured model shown in Fig.˜2. Key geometric, hydrostatic, and hydrodynamic parameters at both model and extrapolated full scales are summarized in Table˜1.
@@
Mean wave height HmHm [m] 2.0366 0.0611
RMS of wave height HrmsHrms [m] 2.2981 0.0689
Encountering frequency (min) $f_{e}$ [Hz] min 0.4000 0.4619
@@
Wave length (min) $\lambda$ [m] min 56.8146 10.7547
Wave length (max) $\lambda$ [m] max 523.7561 14.9574
Non dimensional wave length (min) $\lambda/lpp$ min 1.6881 0.5682
Non dimensional wave length (max) $\lambda/lpp$ max 5.2376 0.9858
Min $H_{m}$ over wave length $Hm/\lambda$ min 0.0358 0.0083
Max $H_{m}$ over wave length $Hm/\lambda$ max 0.0358 0.0358
@@ 2.2.1 Bayesian HDMDc
$\boldsymbol{\sigma_{x}}(t)=\Bigg\{\int_{\lambda^{l}}^{\lambda^{u}}\int_{l_{d_{u}}^{l}}^{l_{d_{u}}^{u}}\int_{l_{d_{x}}^{l}}^{l_{d_{x}}^{u}}\int_{l_{tr}^{l}}^{l_{tr}^{u}}\left[\mathbf{x}(t,l_{tr},l_{d_{x}},\lambda)-\boldsymbol{\mu_{x}}(t)\right]^{2}\\ p(l_{tr})p(l_{d_{x}})p(l_{d_{u}})p(\lambda)dl_{tr}dl_{d_{x}}dl_{d_{u}}d\lambda\Bigg\}^{\frac{1}{2}},$ (15)
@@ 2.3 Performance metrics
The JSD is based on the Kullback-Leibler divergence $D$, given by Eq.˜24, which is the expectation of the logarithmic difference between the probabilities $K$ and $H$, both defined over the domain $\chi$, where the expectation is taken using the probabilities $K$ Kullback1951 The similarity between the distributions is higher when the Jensen-Shannon distance is closer to zero. JSD is upper bounded by $\ln(2)$. When applied to the PDFs evaluated from the bootstrapped time series, the expected value and the quantile function of JSD for $p=0.025$ and $0.975$, defining the lower and upper bound of the 95% confidence interval $U_{\text{JSD}(Q,R)}=\text{JSD}(Q,R)_{q=0.95}-\text{JSD}(Q,R)_{q=0.025}$, are evaluated.

### v2

@@ 1 Test Case
A fiberglass scale model, designated as CNR-INM 2554, was constructed to replicate the Delft design at a scale of 1:33.33, corresponding to a model length ($L_{pp}$) of 3 meters, considering the full-scale prototype to be a notional 100-meter catamaran. An outline of the model’s geometry and body plan is illustrated in Fig.˜1, with a visual of the manufactured model shown in Fig.˜2. Key geometric, hydrostatic, and hydrodynamic parameters at both model and extrapolated full scales are summarized in Table˜1.
@@
Mean wave height $H_{m}$ [m] 2.0366 0.0611
RMS of wave height $H_{rms}$ [m] 2.2981 0.0689
Encountering frequency (min) $f_{e}$ [Hz] min 0.0800 0.4619
@@
Wave length (min) $\lambda_{w}$ [m] min 56.8146 1.7547
Wave length (max) $\lambda_{w}$ [m] max 523.7561 14.9574
Non dimensional wave length (min) $\lambda_{w}/L_{pp}$ min 0.5681 0.5682
Non dimensional wave length (max) $\lambda_{w}/L_{pp}$ max 5.2376 4.9858
Min $H_{m}$ over wave length $H_{m}/\lambda_{w}$ min 0.0039 0.0041
Max $H_{m}$ over wave length $H_{m}/\lambda_{w}$ max 0.0358 0.0358
@@ 2.2.1 Bayesian HDMDc
$\boldsymbol{\sigma_{x}}(t)=\Bigg\{\int_{\lambda^{l}}^{\lambda^{u}}\int_{l_{d_{u}}^{l}}^{l_{d_{u}}^{u}}\int_{l_{d_{x}}^{l}}^{l_{d_{x}}^{u}}\int_{l_{tr}^{l}}^{l_{tr}^{u}}\left[\mathbf{x}(t,l_{tr},l_{d_{x}},l_{d_{u}},\lambda)-\boldsymbol{\mu_{x}}(t)\right]^{2}\\ p(l_{tr})p(l_{d_{x}})p(l_{d_{u}})p(\lambda)dl_{tr}dl_{d_{x}}dl_{d_{u}}d\lambda\Bigg\}^{\frac{1}{2}},$ (15)
@@ 2.3 Performance metrics
The JSD is based on the Kullback-Leibler divergence $D$, given by Eq.˜24, which is the expectation of the logarithmic difference between the probabilities $K$ and $H$, both defined over the domain $\chi$, where the expectation is taken using the probabilities $K$ Kullback1951 The similarity between the distributions is higher when the Jensen-Shannon distance is closer to zero. JSD is upper bounded by $\ln(2)$. When applied to the PDFs evaluated from the bootstrapped time series, the expected value and the quantile function of JSD for $p=0.025$ and $0.975$, defining the lower and upper bound of the 95% confidence interval $U_{\text{JSD}(Q,R)}=\text{JSD}(Q,R)_{q=0.975}-\text{JSD}(Q,R)_{q=0.025}$, are evaluated.